OpenStudy (anonymous):

Factor Completely 14x^2 + 49xy + 21y^2

OpenStudy (anonymous):

7(x+3y)(2x+1y)

OpenStudy (anonymous):

show work ><???

OpenStudy (anonymous):

im on my 3rd math work sheet..

OpenStudy (anonymous):

\[14x^2 + 49xy + 21y^2\]factor out GCF\[=7(2x^2+7xy+3y^2)\]use ac-method if you like (or trial and error)\[=7\left[ 2x^2+6xy+xy+3y^2 \right]\]\[=7\left[ 2x(x+3y)+y(x+3y) \right]\]\[=7(x+3y)(2x+y)\]